OpenStudy (anonymous):

Mr. Trundle used exactly 128 ceramic tiles to cover the floor of a shower. If each tile is a square with 6-inch sides, what is the area of the shower floor?

OpenStudy (ganpat):

Area of each tile = 6 square = 36 sq inch... so total area = 128 * 6...

OpenStudy (anonymous):

768? o:

OpenStudy (ganpat):

sorry, 128 * 36.. Typo ERROR !!!

OpenStudy (ganpat):

4608 sq inch.. !!!!!!!!!

OpenStudy (anonymous):

Oh, okay. So that equals 4806:D

OpenStudy (ganpat):

:D, ^_^

OpenStudy (anonymous):

Oh, you already posted! Can you answer another question?

OpenStudy (ganpat):

and whats that ?? and how abt u give a try once ??

OpenStudy (anonymous):

A bicycle tire has a diameter of 26 inches. If a pebble is stuck in the tire's treads, about how far will the pebble travel with each rotation of the tire? And okay? I guess, but I don't get these questions at all, so give me formulas to plug in my answers. :D

OpenStudy (ganpat):

L = 2* Pi * r .. r = radius of tire..

OpenStudy (anonymous):

So L=2*Pi*26 then solve?

OpenStudy (ganpat):

hold on..

OpenStudy (anonymous):

mk:)

OpenStudy (ganpat):

yes, solve further :-p

OpenStudy (anonymous):

Is the answer 163.28! :D

OpenStudy (ganpat):

u have take dia.. divide it by 2.. so answer would be, 81.64 inch per rotation

OpenStudy (anonymous):

.. I'm still confused .. what is dia????????

OpenStudy (ganpat):

diameter is complete distance of circle.. refer following figure..